OpenStudy (anonymous):

how do you graph this equation 2x + 5y =10

hero (hero):

By finding the x and y intercepts

OpenStudy (anonymous):

ok so i subract 2x from both sides which equals 5y=-2x+10... then do i divide 5 by -2x and 10?

hero (hero):

Basically, in order to plot a line, you need at least two points. Anytime you plot a line, one of the points will go through the x-axis, and the other will go through the y-axis.

OpenStudy (anonymous):

wait how do i do it?

OpenStudy (anonymous):

nvm

hero (hero):

So you will have two points point a = (x,y) point b = (x,y)

hero (hero):

Point a will be on the x-axis Point b will be on the y-axis If point a is on the x-axis, that means y = 0. If point b is on the y-axis that means x = 0.

OpenStudy (anonymous):

ummm i don't think so... i'm pretty sure its something like the way I was doing it but idk

hero (hero):

There's always more than one way. I'm trying to show you the easy way to do it.

OpenStudy (anonymous):

so how do u plug the number in and graph it in this equation form?

hero (hero):

Just let me show you

OpenStudy (anonymous):

ok(:

hero (hero):

So as I was saying: Point a will be on the x-axis Point b will be on the y-axis If point a is on the x-axis, that means y = 0. If point b is on the y-axis that means x = 0.

hero (hero):

So basically, we know the y coordinate of point a and the x coordinate of point b: Point a = (x,0) Point b = (0,y)

hero (hero):

What we have to do now now is find the remaining x and y coordinates of points a and b

hero (hero):

To find the x coordinate of point a, substitute y = 0 into the formula like so: 2x + 5(0) =10

hero (hero):

Now you can solve for x: 2x + 0 = 10 2x = 10 x = 10/2 x = 5

hero (hero):

So that means Point a = (5,0) Are you with me so far?

OpenStudy (anonymous):

yes

hero (hero):

Okay, so now that we've found point a, we can proceed with finding point b. Remember, for point b, we already have the x-coordinate, x = 0, and we need to find the y - coordinate. To find the y-coordinate, substitute x = 0 into the formula and solve for y: 2(0) + 5y = 10 5y = 10 y = 10/2 y = 5 So this means point b = (0,5)

hero (hero):

So now, we have two points point a = (5,0) point b = (0,5) All you have to do now is take some graphing paper, plot both points, take a ruler and draw a line through the points. Don't forget to label the line 2x + 5y = 10. You're done!

hero (hero):

I hope I didn't lose you